OpenStudy (anonymous):

solve for x between 0 to 2pi secx=2cscx

Miracrown (miracrown):

I can guide you through this, but I'll need your help So first of all an always good starting point is to see if we can rewrite both trig functions in terms of other trig functions For example: cotg x can be rewritten as 1/tg x

Miracrown (miracrown):

If we can rewrite cotg x on that way, how can we rewrite sec x on a similar way?

OpenStudy (anonymous):

1/cosx=2(1/sinx)

Miracrown (miracrown):

To review: both sec x and csc x can be rewritten as 1/ another trig function and those are ...

Miracrown (miracrown):

Yes, ur right

OpenStudy (anonymous):

what is the next step?

Miracrown (miracrown):

And is there any other trig function that can be written in terms of both sin x and cos x?

OpenStudy (anonymous):

sinx/cosx=tanx

Miracrown (miracrown):

Yerp! And is there any way we can pit together sin x and cois x at our equation in order to get tan x?

OpenStudy (anonymous):

sinx/cosx=2

OpenStudy (anonymous):

tanx=2

Miracrown (miracrown):

Nice :)

Miracrown (miracrown):

And what would you do next?

OpenStudy (anonymous):

arctan of 2

Miracrown (miracrown):

Mhm!

OpenStudy (anonymous):

1.107

OpenStudy (anonymous):

Is it right?

Miracrown (miracrown):

Yush!!!

OpenStudy (anonymous):

so this is the answet?

Miracrown (miracrown):

Its right, but I think that there may still be other answers missing if we have to think on tan x... so if we have to think on tan x as a function what do you remember about it?

OpenStudy (anonymous):

I have no clue

Miracrown (miracrown):

as you may remember the 3 main trig functions (sin x, cos x and tan x) are all periodic functions that is, the values each of them takes repeat every certain number of units for example, for sin x values repeat every 2pi units and therefore sin(pi)=sin(3pi)

OpenStudy (anonymous):

tanx repeats every pi

Miracrown (miracrown):

And the same happenbs with cos x every 2pi units too but tan x is a little but different since it has some holes on it (I mean, we can't calculate tan(pi/2, for example) and its values get repeated every pi units so then, for example tan(pi/4)=tan(5pi/4)

OpenStudy (anonymous):

4.249

Miracrown (miracrown):

So then, if we know that for x=1.107 radians tan x=2 which other value between 0 and 2pi can happen the same for?

Miracrown (miracrown):

Yes, 4.249 is correct. And is there any other possible value? between 0 and 2pi?

OpenStudy (anonymous):

can I subtract pi too

OpenStudy (anonymous):

1.107-pi

Miracrown (miracrown):

Yes, you can and what would you get by doing so?

OpenStudy (anonymous):

-2.034

Miracrown (miracrown):

right and is that between 0 and 2pi?

OpenStudy (anonymous):

no.It is less than 0

OpenStudy (anonymous):

can I add one more pi

OpenStudy (anonymous):

4.249+pi

Miracrown (miracrown):

And what does that equal to?